# Best Practices for Multi-Fidelity Bayesian Optimization in Materials and Molecular Research

### v2

@@ Methods
Recent work in MFBO for materials discovery has seen the emergence of several metrics to evaluate method performance. To measure the progress of the optimization campaign, regret15, 16, Active Learning Metrics (ALM)19 and campaign efficiency20 have been proposed. In the ML literature, regret is the most common metric, although its specific implementation for MFBO cases is not defined21, 11, 23, 18. To measure absolute MFBO performance, that is, how a specific MFBO run compares to its SFBO counterpart under specific experimental conditions, Acceleration Factor (AF) and Enhancement Factor19, cost difference to discover materials in the 99th best percentile15 and MF advantage20 have been used. Importantly, MFBO performance is also dependent on the available budget, and this absolute metric has to somehow capture this dependency. Only some works have explicitly mentioned 20, 16 or studied15 how low-fidelity informativeness affects the final result, quantifying it using the $R^{2}$ or correlation coefficient between the HF and LF source, respectively. Last, some recent works highlighted the need of better metrics to assess MFBO performance 22, 27. To solve this lack of standardization, and building on the previous works, we propose two key metrics - MFBO regret (r) and discount ($\Delta$) to study MFBO performance under different scenarios. We also define $\rho$ and $R^{2}$ as the metrics to characterize the LF source.
@@
Figure 2: Standardized metrics to compare MFBO and SFBO. Simple regret traces are computed using algorithm 1 and provide a direct comparison between the optimization progress in each method by computing the equivalent number of steps from SFBO in the MFBO trace. Discount ($\Delta$) reflects the savings in optimization cost by computing the normalized difference in budget spent by each algorithm ($\texttt{b}^{\rm sf}(\tilde{r}^{*})$ and $\texttt{b}^{\rm mf}(\tilde{r}^{*})$) to find $\tilde{r}^{*}$ (horizontal arrow, see equation 1). The plot is zoomed in to focus on the $\Delta$ score.
We define a standardized simple regret and a discount metric to quantify the performance of MFBO over SFBO along the optimization. This metric aims to inform how the MFBO optimization progresses with respect to a known reference single-fidelity source. This approach is similar to the idea of ALM introduced in 19 (note that it is an "after-the-fact" metric, meaning that the best candidate must be known beforehand, although this is common to all the benchmarks). Let $f$ be the HF target problem. At a given step $t$ of the SFBO run, the simple regret is defined as $r_{t}=f^{*}-f^{*}_{t}$, where $f^{*}$ is the known global optimum and $f^{*}_{t}$ is the best value of $f$ found so far. The simple regret is computed only at the highest fidelity level in the multi-fidelity runs. To compare multi-fidelity and single-fidelity runs it is necessary to find the corresponding high-fidelity simple regret in the multi-fidelity setting, which can query both the low-fidelity or high-fidelity at a given step. We define a standard algorithm to identify the corresponding HF simple regret at a given budget of an MFBO campaign. The algorithm aligns the ${r}$ values from the MFBO to a common cost scale given by the SFBO cost steps. This approach makes it possible to compare the performance of the MFBO method at each step of the associated SFBO run. Algorithm 1 explains in detail the procedure to compute simple regret both for single-fidelity and multi-fidelity runs. Even if in our work we only evaluate two sources of fidelity, the advantage of this regret computation is that it allows to incorporate as many discrete LF sources as the user wants. A potential disadvantage may be the need of a reference optimization case (the SFBO run in this case), but we consider this acceptable as the final objective of the study is to decide when MFBO is better than SFBO.
@@
We also propose a discount $\Delta$ metric to estimate the savings provided by the MFBO method compared to the SFBO (see figure 2 for a graphical explanation of this metric). This metric is inspired is the Acceleration Factor (AF) previously proposed19. $\Delta$ reflects the difference in budgets required to reach a reference regret value between the multi-fidelity and single-fidelity methods. Specifically, it is computed over two individual SF and MF runs as follows:
@@
$\tilde{r}^{*}$ is the corrected best single-fidelity regret, and $r^{\rm sf}_{\rm max}$ and $r^{\rm sf}_{\rm min}$ are the maximum and minimum regrets in the SFBO run, and $\tau$ is a factor that quantifies the amount of total reduced regret that the user wants to sacrifice in order to get a good MFBO performance. The $\tau$ correction accounts for situations where the MFBO method gets to low regrets faster than the SFBO at the beginning of the optimization, but its performance plateaus in the long term as noted in22. This correction also reflects the dependence of MFBO performance on the available budget, as mentioned in previous work20. We report discounts with $\tau$ = 0.9, although we also investigate the discount obtained for several values of $\tau$ in the synthetic functions benchmark (see SI section A.4. A positive $\Delta$ indicates cost savings with MFBO, meaning it achieves the minimum regret with a lower budget than SFBO. Conversely, a negative $\Delta$ implies that MFBO was more expensive than SFBO. If the MFBO method is unable to reach $\tilde{r}^{*}$, $\Delta$ is set to -1. Note that equation 1 compares individual runs started at the same seed as displayed in Figure 2, and we report the average $\Delta$ over all the seeds onwards.
@@
Synthetic functions are simulated black-box problems where a mathematical expression is used to generate points from a surface to be optimized. In the multi-fidelity case, there are lower accuracy approximations of the target function at a lower cost. These can either have a fixed expression or be generated by biasing the original function using a given parameter $\alpha$. In our case, we use the Branin23 function previously used in MFBO works. We also modify a Park function used in MFBO28 by incorporating a parameter $\alpha$ to modulate the bias with a similar behaviour to the Branin case. The expressions of the functions can be found in the Supplementary Information A.2. The cost of querying the high-fidelity level is set to a value of 1, and the cost of the low-fidelity approximations is set to a fraction of this value.
@@
We use or adapt previously reported benchmarks for real experimental optimization problems in the chemistry domain. The Covalent Organic Frameworks (COFs) dataset was used in a previous work on MFBO16, and consists of 608 candidate COFs encoded in a 14-dimensional vector accounting for their composition and crystal structure. The high-fidelity simulation uses a Markov chain Monte Carlo simulation to compute the adsorption of Kr and Xe in the material, and the associated Kr/Xe selectivity. This simulation is expensive, with an average running time of 230 minutes. In this case, a low-fidelity approximation of the selectivity can be obtained using Henry’s law, reducing the computing time to 15 minutes, giving a $\rho$ of 0.065 for this problem.
@@
The second benchmark is extracted from the FreeSolv library29. It comprises 641 molecules, encoded using RDKit 2D-descriptors and reduced to a 10-D vector using PCA. The high-fidelity is the experimental free solvation energy, whereas the low-fidelity is the computed solvation energy using molecular dynamics (MD). $\rho$ is set up to 0.1 in this case, based on an estimation of the difference between running a solvation measurement and an MD simulation.
@@
The last benchmark is derived from the Alexandria library 30, and it was used in previous work on MFBO for materials discovery15. It comprises 1134 molecules encoded using RDKit 2D-descriptors 31 and reduced to a 10-D vector using PCA. The high-fidelity is the experimental polarizability, whereas the low-fidelity is the computed polarizability at the Hartree-Fock 6-31G+ level of theory. $\rho$ is set up to 0.167 in this case following the previous study.
Results
Shedding light into MFBO failure modes
We test the performance of MFBO methods over their SFBO counterparts on two different synthetic functions, Branin and Park (with 2 and 4 dimensions respectively), on two different scenarios (favorable and unfavorable). Synthetic functions are commonly employed as black-box problems to test the performance of BO algorithms (see Synthetic functions). In the case of MFBO, the output of the functions can be biased using a parameter $\alpha$ to provide lower accuracy information sources. This parameter is related to the informativeness of the LF source (in our case, we quantify the informativeness of the LF by computing the $R^{2}$ with respect to the HF, see Metrics). For the favorable scenario, we tune the $\alpha$ parameter to provide a highly informative LF source (LF $R^{2}>0.9$). We set the cost of querying a HF point to 1 and the cost of a LF point to 0.1 ($\rho$ = 0.1). Figure 3 shows the results for the favorable scenario. In this case, both MFBO methods outperform their SFBO counterparts, with maximum discounts of 0.57 and 0.38 for Branin and Park respectively. The MFBO runs get to lower regrets quicker, spending less resources by exploiting the use of LF and HF sources of information. The resulting optimization runs are therefore desirable as they effectively leverage the access to the LF to guide the optimization of the HF level.
Although under the previous settings MFBO provides a higher performance than SFBO, a change in the LF conditions dramatically affects the result of the optimization. In the unfavorable scenario, we set the cost of the LF source to 0.5 (only half of the HF cost, $\rho$ = 0.5) and decrease the informativeness of the functions ($R^{2}$ $<$ 0.75). In this case, the performance of both methods decreases with respect to the favorable scenario. In Branin, the maximum $\Delta$ drops to 0.23, whereas in Park the trend is reversed and the MFBO loses its advantage over SFBO, with a maximum $\Delta$ of -0.44. These results illustrate how MFBO performance is affected by the cost and the informativeness of the LF source, losing its advantage over SFBO if the LF source is not informative and cheap enough. It also reflects how the MFBO setting is problem-dependent and it may be more robust towards problem changing situation
Apart from problem conditions, model choice may also affect MFBO performance. Previous studies have proposed and used many MF models, mainly GPs refs but also DNNs10, 32. We compared the standard BoTorch multi-fidelity model with a MultiTask GP9 to see how model choice can affect the results. The main difference between the two models lies in how fidelities are encoded in the GP. SI section A shows a preliminary study on the fidelity kernel and how both models offer similar performances in regression and MFBO tasks with the Branin function (figures 7 and 8 respectively). Therefore, we stick to the use of the default model provided in BOTorch, focusing on the external factors that can affect the experiments.
Figure 3: MFBO results are dependent on problem conditions. In the favorable scenario (highly informative and cheap LF source, $\rho$ = 0.1, $R^{2}$ > 0.9), MFBO offers better performance than SFBO on the synthetic functions (maximum $\Delta$ of 0.57 and 0.38 for Branin and Park, respectively). The optimization traces show how lower regrets are found at lower budgets under this scenario. When the conditions are changed and a low informative and more expensive LF is used ($\rho$ = 0.5, $R^{2}$ < 0.75), MFBO performance is degraded. Under this scenario, the optimization traces are similar, and MFBO performance is notably reduced (minimum $\Delta$ of 0.23 and -0.44 for Branin and Park, respectively).
Finding suitable scenarios for MFBO
We investigate in detail the effect of the LF cost and informativeness on MFBO performance (measured by $\Delta$, see Metrics) for the two previous synthetic functions. We run the MFBO method with different combinations of LF cost and informativeness and compare it to SFBO. Figure 4 shows the computed $\Delta$ for each run as a result of the previous parameters. The $x$ axis of the heatmap represent the $\alpha$ value that was used to bias the synthetic function, and the upper plot shows the computed $R^{2}$ for each value. In both cases, a gradient can be observed where the progression towards cheaper and more informative LF sources provides better MFBO performances (higher $\Delta$). In the case of Branin, the lower values of $R^{2}$ in the synthetic functions are associated with a less informative LF approximation, which translates to lower values of $\Delta$. In the same way, in Park $\Delta$ is also decreased when the $R^{2}$ is lower, giving negative results when its value is close to 0. This result is consistent as a more inaccurate source of information is likely to reduce the performance of the MFBO method. In terms of cost, there is also an inverse correlation in both cases, where lower $\rho$ (cheaper LF sources) generate higher discounts. This is an expected trend, as the model can access the approximate sources at a lower cost, allowing a wider exploration of the problem space for the same price.
Although $\Delta$ is reported at a value of $\tau$ = 0.9 for all the cases, we also investigated the effect of $\tau$ on the final discount. In previous works, the MFBO performance was reported to be lower than SFBO at smaller budgets20, and the authors noted that in the long-run MFBO may also lose its advantage over SFBO22. This trend was observed in both functions and AFs, where low values of $\tau$ and a $\tau$ of 1 provided negative discounts (see figure 9 in the SI). Although the temporal dependence of the MFBO is hard to control for the experimenter, due to the uncertainty of when to stop the sampling, the previous figure shows how there exist "sweet spots" where MFBO can exploit the cheap LF source with the available budget to provide an advantage over SFBO.
The results follow the same trend in both cases independently of the acquisition function. Although there are differences in the absolute $\Delta$ values between the two problems (due to the specific landscape of each synthetic function), the regions of high performance are localized in the same areas of the heatmaps. These results provide a qualitative understanding of the effect of cost and informativeness on the performance of MFBO. The general high-performance region is located around cheap and highly informative LF sources, and a good experimental setup for MFBO application should aim to find suitable LF sources. In general, a user should first verify that the cost of the auxiliary experiment is cheap enough to run MFBO. In the same way, they should check and estimate the degree of informativeness of the auxiliary experiment to decide if it provides a reasonable approximation of the real problem (in this case, we propose $R^{2}$ as a guiding metric, but other informativeness measurements could be employed). Figure 5 displays a guiding flowchart following the previous considerations. Using these conditions does not guarantee superior performance over SFBO, but provides a guiding principle for the successful application of MFBO in real scenarios.
Should I use MFBO? $\texttt{Is}\;\rho<\tau_{1}\ =0.2\texttt{?}$ $\texttt{Is}\;R^{2}>\tau_{2}=0.75\texttt{?}$ run SFBO run SFBO run MFBO yesnonoyes
Figure 5: Proposed guidelines to run MFBO. Evaluation of the informativeness and cost of the LF approximation provides a reasonable estimation of the potential success of the MFBO approach over standard SFBO. The flowchart illustrates a potential analysis of the scenario to decide if running MFBO or not by using a threshold of $\rho$ ($\tau_{1}$) and $R^{2}$ ($\tau_{2}$). If the LF source is cheap enough ($\rho<\tau_{1}$) and informative enough ($R^{2}>\tau_{2}$), we recommend running MFBO. We propose $\tau_{1}=0.2$ and $\tau_{2}=0.75$ as a soft threshold for this study based on our empirical investigation.
Application of MFBO guidelines to chemistry problems
We translate the previous recommendations to three benchmarks in chemistry and materials science to showcase their utility in real experimental situations. The tasks correspond to scenarios where a cheap and informative LF source is available under the proposed guidelines (see Methods for a detailed explanation of each benchmark). Figure 6a illustrates the results of MFBO and SFBO in these scenarios. In all cases, the MFBO method can get to lower regrets at a lower cost than the SFBO by leveraging the cheaper source of information at hand. In the COFs benchmark (problem settings: $\rho$ = 0.065, $R^{2}$ = 0.98), the MFBO gets to the COF with the highest computed Kr/Xe selectivity in the allocated budget, whereas the SFBO does not find it. In the polarizability benchmark (problem settings: $\rho$ = 0.167, $R^{2}$ = 0.99), the MFBO provides a maximum discount of 0.6 over the SFBO. In this case, the use of a cheap Hartree-Fock computation is successfully combined with the experimental measurements to find the molecule with the highest polarizability. In the solvation energy benchmark (problem settings: $\rho$ = 0.1, $R^{2}$ = 0.88), the MFBO finds the molecule with the highest solvation energy but the SFBO not, as in the COFs case. This example also integrates computed values with experimental measurements to reduce the spent budget. Figure 6b shows the proportion of queries to each fidelity level in the different tasks. By distributing the available budget between the HF and LF levels, the MFBO methods can integrate the cheaper information of the LF approximations to improve the optimization results. The number of calls to each query depends on the specific acquisition function and task. However, in all cases the proportion of calls to the HF level is lower than 0.4, suggesting that this proportion can provide a threshold for a successful MFBO optimization. A previous study suggested a proportion of LF-HF points of 5:1 as the optimal case for good MFBO performance20, as is the case for the COFs and polarizability benchmarks. In these real scenarios, the MFBO methods can exploit the low accuracy but cheap LF levels by querying them more often than the HF level.
Figure 6: Molecules and materials benchmark results. a) Optimization performance for the MFBO and the SFBO methods in each benchmark (see Methods for a description of each task). The maximum $\Delta$ in each case is 0.68, 0.45 and 0.56 for the COFs, solvation energy and polarizabiliry tasks, respectively. In all the cases, MFBO is able to achieve lower regrets than SFBO by spending less budget. b) Proportion of queries to each fidelity source in the MFBO setting. The MFBO method balances information gain and budget spent by querying more often the LF values (between 0.6 and 0.8 in the ratio of queries) while also using information from HF values. c) Experimental negative example. Molecule polarizability benchmark conditions were changed so the LF source was expensive and less informative. The resulting $\Delta$ is -0.07 and 0.13 for the MES and EI families respectively, showing how unfavorable conditions can override the MFBO advantage over SFBO in a real experimental setting.
We also include a negative example where the conditions of one of these benchmarks are changed to an unfavorable scenario and the associated original MFBO performance is overridden. We artificially change the LF conditions of the molecule polarizability benchmark ($\rho$ = 0.5 and $R^{2}$ = 0.49) to illustrate how $\Delta$ is reduced with respect to the previous case. Figure 6c shows the result for the optimization, where $\Delta$ is reduced from a maximum value of 0.42 to 0.13 in the EI case (for MES, the performance decreases from 0.21 to -0.07), providing almost no advantage over SFBO if compared with the previous favorable scenario (similar to the results described for the synthetic functions, see Results). The previous benchmarks show how the application of MFBO in suitable scenarios provides an advantage over standard SFBO cases and effectively decreases the optimization cost. By locating experiments where the cost and informativeness of the LF source are adequate, we can successfully translate the advantages of MFBO into a realistic experimental task.
Discussion
Multi-fidelity Bayesian optimization (MFBO) offers a promising approach to reducing costs in optimization by leveraging inexpensive, approximate sources of information. In the chemical sciences, MFBO can accelerate optimization while maintaining cost efficiency. However, its utility is not universal; in some cases, it may prove less effective or even counterproductive23. To address these concerns, we conducted a comprehensive study to determine under which conditions MFBO outperforms standard single-fidelity Bayesian optimization (SFBO).
Building on previous work in the field, we introduce two key metrics for comparison: a standardized regret, $r$, and a discount metric, $\Delta$ (see Metrics). Our analysis of synthetic benchmarks demonstrates that MFBO performance is highly dependent on the characteristics of the low-fidelity (LF) source. Specifically, not all LF conditions are conducive to the effective use of MFBO. To systematically identify favorable conditions, we performed a grid-based exploration of LF cost and informativeness against the discount metric, $\Delta$. The results indicate that both $\rho$ and $R^{2}$ significantly influence the discount, with the most favorable outcomes achieved when the LF source is both highly informative and low-cost. Based on these findings, we developed a flowchart to guide the application of MFBO (see Figure 5). These guidelines were validated using three real-world experimental benchmarks in molecular and materials discovery, demonstrating that MFBO consistently outperforms SFBO in reducing optimization costs, provided the LF source is sufficiently inexpensive and informative.
Our work provides a structured decision-making framework for determining the applicability of MFBO. While our analysis focused on a specific surrogate model and two acquisition functions, the approach is extensible to other surrogate models and acquisition function families. Future research will expand this framework to include models such as Bayesian neural networks (BNNs)33, and explore additional experimental applications, evaluating the impact of varied feature spaces or additional fidelity levels. This study offers valuable insights for practitioners seeking to integrate MFBO into their experimental workflows, paving the way for more routine application of MFBO in chemical and materials optimization.
Code and data availability
@@ Discussion
The COFs benchmark data is available at https://github.com/SimonEnsemble/multi-fidelity-BO-of-COFs-for-Xe-Kr-seps., and it was extracted from 16. The polarizability dataset is available at https://zenodo.org/records/100471130. It was used in previous MFBO work15. The solvation energy dataset is extracted from 29. The code is available in this Github repository: https://github.com/atinary-technologies/chem-MFBO.git. It includes the instructions to run all the experiments described in this work.

### v3

@@ Methods
Recent work in MFBO for materials discovery has seen the emergence of several metrics to evaluate method performance. To measure the progress of the optimization campaign, regret18, 19, Active Learning Metrics (ALM)22 and campaign efficiency23 have been proposed. In the ML literature, regret is the most common metric, although its specific implementation for MFBO cases is not defined26, 14, 29, 21. To measure absolute MFBO performance, that is, how a specific MFBO run compares to its SFBO counterpart under specific experimental conditions, Acceleration Factor (AF) and Enhancement Factor22, cost difference to discover materials in the 99th best percentile18 and MF advantage23 have been used. Importantly, MFBO performance is also dependent on the available budget, and this absolute metric has to somehow capture this dependency. Only some works have explicitly mentioned 23, 19 or studied18 how low-fidelity informativeness affects the final result, quantifying it using the $R^{2}$ or correlation coefficient between the HF and LF source, respectively. Last, some recent works highlighted the need of better metrics to assess MFBO performance 27, 28. To solve this lack of standardization, and building on the previous works, we propose two key metrics - MFBO regret (r) and discount ($\Delta$) to study MFBO performance under different scenarios. We also define $\rho$ and $R^{2}$ as the metrics to characterize the LF source.
@@
Figure 6: Standardized metrics to compare MFBO and SFBO. Simple regret traces are computed using algorithm 1 and provide a direct comparison between the optimization progress in each method by computing the equivalent number of steps from SFBO in the MFBO trace. Discount ($\Delta$) reflects the savings in optimization cost by computing the normalized difference in budget spent by each algorithm ($\texttt{b}^{\rm sf}(\tilde{r}^{*})$ and $\texttt{b}^{\rm mf}(\tilde{r}^{*})$) to find $\tilde{r}^{*}$ (horizontal arrow, see equation 1). The plot corresponds to the average trace of the MFEI and SFEI methods on an example favorable Branin-2D case (see Figure 3), and it is zoomed in to focus on the $\Delta$ score.
We define a standardized simple regret and a discount metric to quantify the performance of MFBO over SFBO along the optimization. This metric aims to inform how the MFBO optimization progresses with respect to a known reference single-fidelity source. This approach is similar to the idea of ALM introduced in 22 (note that it is an "after-the-fact" metric, meaning that the best candidate must be known beforehand, although this is common to all the benchmarks). Let $f$ be the HF target problem. At a given step $t$ of the SFBO run, the simple regret is defined as $r_{t}=f^{*}-f^{*}_{t}$, where $f^{*}$ is the known global optimum and $f^{*}_{t}$ is the best value of $f$ found so far. The simple regret is computed only at the highest fidelity level in the multi-fidelity runs because this level is the real optimization target and the true reference to measure the method’s performance. To compare multi-fidelity and single-fidelity runs it is necessary to find the corresponding high-fidelity simple regret in the multi-fidelity setting, which can query both the low-fidelity or high-fidelity at a given step. We define a standard algorithm to identify the corresponding HF simple regret at a given budget of an MFBO campaign. The algorithm aligns the ${r}$ values from the MFBO to a common cost scale given by the SFBO cost steps. This approach makes it possible to compare the performance of the MFBO method at each step of the associated SFBO run. Algorithm 1 explains in detail the procedure to compute simple regret both for single-fidelity and multi-fidelity runs. Even if in our work we only evaluate two sources of fidelity, the advantage of this regret computation is that it allows to incorporate as many discrete LF sources as the user wants. A potential disadvantage may be the need of a reference optimization case (the SFBO run in this case), but we consider this acceptable as the final objective of the study is to decide when MFBO is better than SFBO.
@@
We also propose a discount $\Delta$ metric to estimate the savings provided by the MFBO method compared to the SFBO (see figure 6 for a graphical explanation of this metric). This metric is inspired by the Acceleration Factor (AF) and Active Learning Metrics (ALM) previously proposed and used in MFBO work22. $\Delta$ reflects the difference in budgets required to reach a reference regret value between the multi-fidelity and single-fidelity methods. This metric can be viewed as a combination of the simple regret and the total spent budget, reflecting how MFBO results are an interplay of the HF value optimization and the total budget spent. Specifically, it is computed over two individual SF and MF runs as follows:
@@
$\tilde{r}^{*}$ is the corrected best single-fidelity regret, and $r^{\rm sf}_{\rm max}$ and $r^{\rm sf}_{\rm min}$ are the maximum and minimum regrets in the SFBO run, and $\tau$ is a factor that quantifies the amount of total reduced regret that the user wants to sacrifice in order to get a good MFBO performance. The $\tau$ correction accounts for situations where the MFBO method gets to low regrets faster than the SFBO at the beginning of the optimization, but its performance plateaus in the long term as noted in27. This correction also reflects the dependence of MFBO performance on the available budget, as mentioned in previous work23.$\tau$ refers to the slack a user will give with respect to the converging regret of SFBO, or how good the user wants their output to be with respect to the minimum regret achievable by SFBO.
We report discounts with $\tau$ = 0.9, although we also investigate the discount obtained for several values of $\tau$ in the synthetic functions benchmark (see SI section A.4. A positive $\Delta$ indicates cost savings with MFBO, meaning it achieves the minimum regret with a lower budget than SFBO. Conversely, a negative $\Delta$ implies that MFBO was more expensive than SFBO. If the MFBO method is unable to reach $\tilde{r}^{*}$, $\Delta$ is set to -1. Note that equation 1 compares individual runs started at the same seed as displayed in Figure 6, and we report the average $\Delta$ over all the seeds onwards.
@@
Synthetic functions are simulated black-box problems where a mathematical expression is used to generate points from a surface to be optimized. In the multi-fidelity case, there are lower accuracy approximations of the target function at a lower cost. These can either have a fixed expression or be generated by biasing the original function using a given parameter $\alpha$. In our case, we use the Branin29 function previously used in MFBO works. We also modify a Park function used in MFBO38 by incorporating a parameter $\alpha$ to modulate the bias with a similar behaviour to the Branin case. The expressions of the functions can be found in the Supplementary Information A.2, and a visualization of the $\alpha$ parameter’s effect on the Branin function (figure 9). The cost of querying the high-fidelity level is set to a value of 1, and the cost of the low-fidelity approximations is set to a fraction of this value.
@@
We use or adapt previously reported benchmarks for real experimental optimization problems in the chemistry domain. The Covalent Organic Frameworks (COFs) dataset was used in a previous work on MFBO19, and consists of 608 candidate COFs encoded in a 14-dimensional vector accounting for their composition and crystal structure. The high-fidelity simulation uses a grand canonical Markov chain Monte Carlo simulation to compute the adsorption of Kr and Xe in the material, and the associated Xe/Kr selectivity. This simulation is expensive, with an average running time of 230 minutes. In this case, a low-fidelity approximation of the selectivity can be obtained using Henry’s law with a classical force field, reducing the computing time to 15 minutes, giving a $\rho$ of 0.065 for this problem.
@@
The second benchmark is extracted from the FreeSolv library39. It comprises 641 molecules, encoded using RDKit 2D-descriptors and reduced to a 10-D vector using PCA. The high-fidelity is the experimental free solvation energy, whereas the low-fidelity is the computed solvation energy using molecular dynamics (MD). $\rho$ is set up to 0.1 in this case, based on an estimation of the difference between running a solvation measurement and an MD simulation.
@@
The last benchmark is derived from the Alexandria library 40, and it was used in previous work on MFBO for materials discovery18. It comprises 1134 molecules encoded using RDKit 2D-descriptors 41 and reduced to a 10-D vector using PCA. The high-fidelity is the experimental polarizability, whereas the low-fidelity is the computed polarizability at the Hartree-Fock 6-31G+ level of theory. $\rho$ is set up to 0.167 in this case following the previous study.
Data availability
@@
The COFs benchmark data is available at https://github.com/SimonEnsemble/multi-fidelity-BO-of-COFs-for-Xe-Kr-seps., and it was extracted from 19. The polarizability dataset is available at https://zenodo.org/records/100471140. It was used in previous MFBO work18. The solvation energy dataset is available at https://escholarship.org/uc/item/6sd403pz and is based on 39.
Code availability
The code is available in this Github repository https://github.com/atinary-technologies/chem-MFBO.git (ref. 42). It includes the instructions to run all the experiments described in this work.